# Bayesian Hierarchical Probabilistic Forecasting of Intraday Electricity Prices

### v1

@@ 3.3 Construction of feature space for regression
Regressor
Description
DA market and live CID market
@@
$y-P_{\mathrm{da}}$
Price spread between live IDFull and DA price.
$V_{\mathrm{cid}}-V_{\mathrm{da}}$
Volume spread between live CID volume and DA volume.
$\eta_{\mathrm{cid}}(\Delta p)-\eta_{\mathrm{da}}(\Delta p)$
@@
Energy production forecasts in reference to DA market and live CID market
$E_{\mathrm{res}}=E_{\mathrm{tot}}-E_{\mathrm{sol}}-E_{\mathrm{won}}-E_{\mathrm{woff}}$
Residual energy production
$E_{\mathrm{tot}}-V_{\mathrm{da}}$
DA excess energy production.
$E_{\mathrm{res}}-V_{\mathrm{da}}$
DA excess residual energy production.
$E_{\mathrm{tot}}-V_{\mathrm{cid}}$
CID excess energy production.
$E_{\mathrm{res}}-V_{\mathrm{cid}}$
CID excess residual energy production.
Energy consumption forecasts in reference to DA market and live CID
$C_{\mathrm{res}}=E_{\mathrm{cons}}-E_{\mathrm{sol}}-E_{\mathrm{won}}-E_{\mathrm{woff}}$
Residual energy consumption
$E_{\mathrm{cons}}-V_{\mathrm{da}}$
DA excess energy consumption.
$C_{\mathrm{res}}-V_{\mathrm{da}}$
DA excess residual energy consumption.
$E_{\mathrm{cons}}-V_{\mathrm{cid}}$
CID excess energy consumption.
$C_{\mathrm{res}}-V_{\mathrm{cid}}$
CID excess residual energy consumption.
Energy production forecast shifts
@@
$E_{\mathrm{won}}^{\mathrm{id}}\!+\!E_{\mathrm{woff}}^{\mathrm{id}}-E_{\mathrm{won}}^{\mathrm{da}}\!-\!E_{\mathrm{woff}}^{\mathrm{da}}$
Difference in intraday and day-ahead wind energy forecast.
@@
Differences to time-lagged values of all features
$\Delta_{\mathrm{h}}(x)=x_{(d,h)}(\tau)-x_{(d,h)-1\mathrm{h}}(\tau)$
Maintaining forecast creation time $\tau$, the difference of all considered features $x_{(d,h)}$ for delivery hour $h$ on day $d$ to the value of that feature for the previous delivery hour is added to the feature space.
$\Delta_{\mathrm{d}}(x)=x_{(d,h)}(\tau)-x_{(d,h)-1\mathrm{d}}(\tau\!-\!1\mathrm{d})$
Changing the day of the forecast creation time $\tau$ to the previous day, the difference of all considered features $x_{(d,h)}$ for delivery hour $h$ on day $d$ to the value of that feature for the same delivery hour $h$ on the previous day $d-1$ is added to the feature space.
Table 3: Summary of all features calculated from base regressors listed in Table 1. The regressors from Tables 1 and 2 together with the variables defined here span the feature space for the forecast model.
@@
We organise the feature values in a design matrix $X(d,h,\tau)$, i.e. each row constitutes a data point containing all feature values for a specific day previous to delivery day $d$ we are forecasting. For all data points, we fix the delivery hour $h$, as mentioned at the beginning of the section. All information that would only become available later than the forecast creation time $\tau$ is truncated, mainly pertaining live CID values and power forecasts. To ensure proper learning from the historical data, $\tau$ is set with respect to day $d$, and for each past $d-\Delta d$ also shifted to this day, such that we get for $n+1$ data points
$X(d,h,\tau)=\big{(}$ $\,x(d\!-\!n,\,h,\,\tau\!-\!n\mathrm{d}),\;x(d\!-\!n\!+\!1,\,h,\,\tau\!-\!n\mathrm{d}\!+\!1),$
$\;\dots,\;x(d,\,h,\,\tau)\,\big{)}^{\mathrm{T}},$ (1)
where $x(d,\,h,\,\tau)$ is a feature vector representing one data point pertaining delivery hour $h$ on delivery day $d$ with information available at creation time $\tau$. With $\tau-n\mathrm{d}$ we denoted shifting the creation time by $n$ days to the past, preserving the time information.
@@
$y(d,h)=\big{(}$ $\,\mathrm{eod}_{d-n}(P_{\mathrm{idfull}}),\;\mathrm{eod}_{d-n+1}(P_{\mathrm{idfull}}),$
@@
with weight vector $w=(w_{1},\dots,w_{m})^{\mathrm{T}}$ and normal random variable $\epsilon\sim\mathcal{N}(0,\sigma^{2})^{m}$, and where we omitted the dependencies.
@@ 3.4 Feature selection
In view of a total feature space dimension of $m=351$, and less than $n=730$ data points available per forecast, feature selection becomes mandatory. An additional difficulty arises from high co-linearities between the features. Reasons for these co-linearities include correlations between markets, common causes for increasing and decreasing energy production and consumption, market participants acting on common ground, and features calculated from base regressors.
@@
is added to the least-square regression of (3) with the choice $p=1$. The strength of the penalty is controlled via a Lagrange parameter $\lambda$, which often is set via cross-validation. With increasing $\lambda$, more weights are shrunk to zero, thus performing effective feature selection. However, in case of strong co-linearities, the LASSO tends to choose features that do not generalise well Su et al. (2017).
Instead, we propose Orthogonal Matching Pursuit (OMP) (Pati et al., 1993) for the task of feature selection. Matching Pursuit is a greedy algorithm that approximates the solution of the sparse regression problem corresponding to a $\ell_{0}$-regularisation via a stepwise iteration through feature space. The $\|w\|_{0}$ penalty imposes a constraint on the minimisation of the number of non-zero weights $w_{i}$ to the least-square regression, which has been shown to be superior over the LASSO and other regularisation techniques (Tropp and Gilbert, 2007; Hastie et al., 2020), but belongs to the NP-hard complexity class. The orthogonal extension, OMP, additionally removes from the target variable the orthogonal projection of the feature selected to the target in each iteration step. This procedure improves convergence and provides additional robustness against co-linearities. We use the algorithm implemented in the python package scikit-learn (Rubinstein et al., 2008; Pedregosa et al., 2011).
Before and after feature selection, both the features in $X$ and the target vector $y$ are standardised to zero mean and unit variance. In addition, zero-valued feature vectors (e.g. solar power at night) and those with missing values (e.g. live $P_{\mathrm{id3}}$ for $\tau<h-3$) are removed.
@@ 4 The forecast model
Based on the data model and feature selection described in the previous section, outputting design matrix $X=X(d,h,\tau)$ and target vector $y=y(d,h)$, we now introduce the Bayesian forecast model used in this study. To this end, we separate historical data $(\bar{y},\bar{X})$, comprising the $n$ data points for days $d\!-\!n,\dots,d\!-\!1$, from the most recent data point $(y^{\ast},X^{\ast})$ for delivery day $d$. Note that at this stage $(\bar{y},\bar{X})$ and $(y^{\ast},X^{\ast})$ have already undergone feature selection.
As probabilistic model for the random target variable $Y$, we extract a Gaussian likelihood from the linear model (3),
@@
For the weight vector $w$ we choose the normal distribution as prior and independently the Gamma distribution for the standard deviation $\sigma$,
@@
Note that this choice imposes $\ell_{2}$-regularisation on $w$ (Ridge regression), c.f. (4). We also tested a Laplacian prior (equivalent to $\ell_{1}$-regularisation) for $w$ but did not find any notable improvement.
The parameters $\mu_{w}$ and $\sigma_{w}$ of the prior for weights $w$ are determined in an empirical Bayes approach. For the mean we choose the ordinary least-square (OLS) estimate $\mu_{w}=C\,\bar{X}^{\mathrm{T}}\,\bar{y}$ with $C=(\bar{X}^{\mathrm{T}}\,\bar{X})^{-1}$. For the standard deviation we make use of the normality of OLS estimates and set $\sigma_{w}=\frac{S}{n}\,\mathrm{diag}(C)$ with $S=(\bar{y}-\mu_{w}\,\bar{X}^{\mathrm{T}})\,\bar{y}$. Similarly, we set the mode $m_{\sigma}=\frac{\alpha-1}{\beta}$ of the Gamma distribution to $1$ to account for standardised data, i.e. $\alpha=m_{\sigma}\beta+1$. Finally, to still ensure weakly informative priors, we found that $\beta=\frac{1}{2}$ is an adequate choice.
@@
Bayesian point estimates for $w$ and $\sigma$ may now be obtained by maximizing the posterior $p_{\mathrm{post}}(w,\sigma\,|\,\bar{X},\bar{y})$ (MAP estimates), which plugged into the likelihood (5) and replacing $\bar{X}$ by $X^{\ast}$ yields a probabilistic forecast for the target $Y$. However, in doing so, we would discard valuable information about the uncertainty of $w$ and $\sigma$ captured by the posterior distribution. To fully incorporate this uncertainty, we instead determine the posterior predictive distribution (PPD)
@@ 5.1 Probabilistic forecasts
We illustrate our forecast procedure with a few examples. Two individual forecasts are shown in Figure 5. Since a typical reference price for the CID market is the DA price $P_{\mathrm{da}}$, we consider the spread between the IDFull and the DA price, $Y-P_{\mathrm{da}}$. Due to the probabilistic forecast, there are several options to extract a point estimate from the estimated predictive distribution $\hat{p}_{\mathrm{ppd}}(y)$, e.g. the mode, median or mean of $\hat{p}_{\mathrm{ppd}}(y)$. In terms of modes, an additional ambiguity arises when $\hat{p}_{\mathrm{ppd}}(y)$ is multimodal. Once a point estimate is extracted from $\hat{p}_{\mathrm{ppd}}(y)$, credible intervals of various credibility levels $\alpha$ can be extracted in various ways, which shall serve as prediction intervals (PIs) in this study. We found the following procedure most robust and accurate in our forecast study.
Figure 5: Example forecasts for day $d=$ 24 August 2022, $h=0$ on the left and day $d=$ 5 July 2022, hour $h=7$ on the right. The forecasts are represented by estimated posterior predictive distributions $\hat{p}_{\mathrm{ppd}}(y)$, from which credible intervals and point estimates are extracted. The multi-modal example on the right illustrates the procedure to pick a prediction interval and a point estimate using highest density intervals (HDIs). The intervals numbered 1, 2, 3, and 4 are intervals that form as intersections with horizontal cuts, c.f. (11), and these intervals are special in that these intervals would fuse to larger intervals for a slightly lower cut. From these fusing intervals, the interval with the largest credibility to width ratio is chosen, c.f. (12), which in this example is interval 1. This interval serves as the PI $\hat{I}$, and the median within this interval as point estimate $\hat{y}$, c.f. (13).
@@
For intermediate $p_{\mathrm{cut}}$, sub-intervals may develop which fuse again to larger credible intervals as $p_{\mathrm{cut}}$ is decreased. In order to choose a PI from all intervals $[l_{i}(p_{\mathrm{cut}}),u_{i}(p_{\mathrm{cut}})]$, we consider all fusing intervals $\{I_{\mathrm{fuse}}\}$, and of these pick $\hat{I}=[\hat{y}_{l},\hat{y}_{u}]$ that maximizes the credibility over width ratio,
@@
This procedure to determine PIs $\hat{I}$ and point estimates $\hat{y}$ in the general case of multimodal predictive distributions is also illustrated in Figure 5. In Figure 6, we show two examples in which the probabilistic forecasts for all hours of a day are depicted. These full day forecasts demonstrate the typical behaviour that our forecast model, namely identifying the live IDFull as the most important regressor, supporting the weak-form efficiency hypothesis. Nevertheless, the point estimates $\hat{y}$ often do correct the live IDFull values towards the ground truth $y^{\ast}$. We furthermore observe that forecasts with a larger lag between creation time $\tau$ and delivery begin $h$ are characterised by broader prediction intervals, indicating that our forecast model correctly represents forecast uncertainty.
XScenario (a), 7 July 2022 XX Scenario (e), 5 October 2022
Figure 6: Exemplary full day forecasts for all delivery hours $h$ of day $d=\;$7 July 2022 on the left and all delivery hours $h$ of day $d=\;$5 October 2022 on the right. The forecast on the left is part of scenario (a), where the forecast creation time $\tau$ is fixed to 23:00 pm on day $d-1$, implying an increasing forecast lag of $h_{\mathrm{lag}}=h+1$ along the horizontal axis. The forecast on the right is part of scenario (e), where the lag between forecast and delivery begin is fixed to 1 hour, i.e. $h_{\mathrm{lag}}=1$, implying forecast creation times $\tau=h-1$. The vertical colour bars represent the topology of the respective predictive distributions $\hat{p}_{\mathrm{ppd}}(y)$ by means of $\alpha(p_{\mathrm{cut}})$ for 100 values of $p_{\mathrm{cut}}$, c.f. (11).

### v2

@@ 3.3 Construction of feature space for regression
Regressor Description
DA market and live CID market I
@@
$y-P_{\mathrm{da}}$ Price spread between live IDFull and DA price.
$V_{\mathrm{cid}}-V_{\mathrm{da}}$ Volume spread between live CID volume and DA volume.
@@
Energy production forecasts in reference to DA market and live CID market I
$E_{\mathrm{res}}=E_{\mathrm{tot}}-E_{\mathrm{sol}}-E_{\mathrm{won}}-E_{\mathrm{woff}}$ Residual energy production
$E_{\mathrm{tot}}-V_{\mathrm{da}}$ DA excess energy production.
$E_{\mathrm{res}}-V_{\mathrm{da}}$ DA excess residual energy production.
$E_{\mathrm{tot}}-V_{\mathrm{cid}}$ CID excess energy production.
$E_{\mathrm{res}}-V_{\mathrm{cid}}$ CID excess residual energy production.
Energy consumption forecasts in reference to DA market and live CID I
$C_{\mathrm{res}}=E_{\mathrm{cons}}-E_{\mathrm{sol}}-E_{\mathrm{won}}-E_{\mathrm{woff}}$ Residual energy consumption
$E_{\mathrm{cons}}-V_{\mathrm{da}}$ DA excess energy consumption.
$C_{\mathrm{res}}-V_{\mathrm{da}}$ DA excess residual energy consumption.
$E_{\mathrm{cons}}-V_{\mathrm{cid}}$ CID excess energy consumption.
$C_{\mathrm{res}}-V_{\mathrm{cid}}$ CID excess residual energy consumption.
Energy production forecast shifts I
@@
$E_{\mathrm{won}}^{\mathrm{id}}\!+\!E_{\mathrm{woff}}^{\mathrm{id}}-E_{\mathrm{won}}^{\mathrm{da}}\!-\!E_{\mathrm{woff}}^{\mathrm{da}}$ Difference in intraday and day-ahead wind energy forecast.
@@
Differences to time-lagged values of all features I
$\Delta_{\mathrm{h}}(x)=x_{(d,h)}(\tau)-x_{(d,h)-1\mathrm{h}}(\tau)$ Maintaining forecast creation time $\tau$, the difference of all considered features $x_{(d,h)}$ for delivery hour $h$ on day $d$ to the value of that feature for the previous delivery hour is added to the feature space.
$\Delta_{\mathrm{d}}(x)=x_{(d,h)}(\tau)-x_{(d,h)-1\mathrm{d}}(\tau\!-\!1\mathrm{d})$ Changing the day of the forecast creation time $\tau$ to the previous day, the difference of all considered features $x_{(d,h)}$ for delivery hour $h$ on day $d$ to the value of that feature for the same delivery hour $h$ on the previous day $d-1$ is added to the feature space.
Table 3: Summary of all features calculated from base regressors listed in Table 1 and external regressors given in Table 2. The regressors from Tables 1 and 2 combined with the variables defined here span the complete feature space for the forecast model.
@@
We organise the feature values in a $(n+1)\times m$ design matrix $X(d,h,\tau)$, i.e. each row constitutes a data point containing all $m$ feature values for a specific day $d-k$, where $d$ is the delivery day we are forecasting and $k=0,...,n$. For all data points, we fix the delivery hour $h$, as mentioned at the beginning of the section. All information that would only become available later than the forecast creation time $\tau$ is truncated, mainly pertaining live CID values and power forecasts. To ensure proper learning from the historical data, $\tau$ is set with respect to day $d$, and for each past day $d-k$ also shifted to this day, such that we get for $n+1$ data points
$X(d,h,\tau)=\Big{(}$ $\;x\big{(}\,d\!-\!n,\,h,\,\tau\!-\!n\mathrm{d}\,\big{)},\;x\big{(}\,d\!-(\!n\!-\!1),\,h,\,\tau\!-(n\!-\!1)\mathrm{d}\,\big{)},$
$\;\dots,\;x\big{(}d,\,h,\,\tau\big{)}\;\Big{)}^{\mathrm{T}},$ (1)
where $x(d-k,\,h,\,\tau)$ is a feature vector representing one data point pertaining delivery hour $h$ on delivery day $d-k$ with information available at creation time $\tau$. With $\tau-n\mathrm{d}$ we denoted shifting the creation time by $n$ days to the past, preserving the time information.
@@
$y(d,h)=\big{(}$ $\,\mathrm{eod}_{d-n}(P_{\mathrm{idfull}}),\;\mathrm{eod}_{d-(n-1)}(P_{\mathrm{idfull}}),$
@@
with weight vector $w=(w_{1},\dots,w_{m})^{\mathrm{T}}$ and normal random variable $\epsilon\sim\mathcal{N}(0,\sigma^{2})^{m}$, omitting the dependence on $d$, $h$, $\tau$.
@@ 3.4 Feature selection
In view of a total feature space dimension of $m=351$, and less than $n=730$ data points available per forecast, feature selection becomes mandatory. An additional difficulty arises from high collinearities between the features. Reasons for these collinearities include correlations between markets, shared causes for increasing and decreasing energy production and consumption, market participants acting on common ground, and features calculated from base regressors.
@@
is added to the least-square regression of (3) with the choice $p=1$. The strength of the penalty is controlled via a Lagrange parameter $\lambda$. With increasing $\lambda$, more weights are shrunk to zero, thus performing effective feature selection. However, in case of strong collinearities, the LASSO tends to choose features that do not generalise well Su et al. (2017).
Instead, we propose Orthogonal Matching Pursuit (OMP) (Pati et al., 1993) for the task of feature selection. Matching Pursuit is a greedy algorithm that approximates the solution of the sparse regression problem corresponding to a $\ell_{0}$-regularisation via a stepwise iteration through feature space. The $\|w\|_{0}$ penalty imposes a constraint on the minimisation of the number of non-zero weights $w_{i}$ to the least-square regression, which has been shown to be superior over the LASSO and other regularisation techniques (Tropp and Gilbert, 2007; Hastie et al., 2020), but belongs to the NP-hard complexity class. The orthogonal extension, OMP, additionally removes from the target variable the orthogonal projection of the feature selected to the target in each iteration step. This procedure improves convergence and provides additional robustness against collinearities. We used the algorithms implemented as linear_model.OrthogonalMatchingPursuit and linear_model.LassoCV in the python package scikit-learn (Rubinstein et al., 2008; Pedregosa et al., 2011).
The hyperparameter for LASSO is the Lagrange parameter $\lambda$, while for OMP it is the maximum number of features, $n_{\mathrm{feat}}$. Typically, $\lambda$ for LASSO is set via cross-validation (CV). However, for better comparison with OMP, we explored optimising $\lambda$ to meet the $n_{\text{feat}}$ constraint but found that the CV built into LassoCV performed better. The hyperparameters of LassoCV had an insignificant influence on the results; hence, we used the default settings (e.g., 5-fold CV). For OMP, the greedy algorithm terminates when either the loss cannot be further improved or the cut-off $n_{\mathrm{feat}}$ is reached. We use $n_{\mathrm{feat}}=20$, ensuring that the number of features determined by OMP typically stays well below the cut-off, except in a few cases with extreme outliers in the data.
To apply feature selection, the data is first cleaned: data points with a few missing values (e.g., occasional gaps in power forecasts, $<0.1\%$) are eliminated, and features with considerable missing values (e.g., live $P_{\mathrm{id3}}$ for $\tau<h-3$) are removed. Then, both the features in $X$ and the target vector $y$ are standardised to zero mean and unit variance. The selected features by OMP and LASSO are then taken from the original data, the cleaning procedure is repeated, and $X$ and $y$ are standardised again. In this way, most missing value problems resolve themselves after feature selection, reducing the need to remove missing data points to a minimum. Finally, zero-valued feature vectors (e.g., solar power at night) are removed if not already deselected by feature selection.
@@ 4 The forecast model
Based on the data model and feature selection described in the previous section, outputting design matrix $X=X(d,h,\tau)$ and target vector $y=y(d,h)$, we now introduce the Bayesian forecast model used in this study. To this end, we separate historical data $(\bar{X},\bar{y})$, comprising the $n$ data points for days $d\!-\!n,\dots,d\!-\!1$, from the most recent data point $(X^{\ast},y^{\ast})$ for delivery day $d$. Note that at this stage $(\bar{X},\bar{y})$ and $(X^{\ast},y^{\ast})$ have already undergone feature selection.
As a probabilistic model for the random target variable $Y$, we extract a Gaussian likelihood from the linear model (3),
@@
Note that any distribution could be chosen here, but without indication of a specific uncertainty structure, we apply the principle that the simplest explanation is usually the best. A Gaussian likelihood also fits well with the empirical Bayes approach explained below. The generality of the model lies in the fact that the posterior predictive distribution is a compound distribution between the likelihood and the posterior, which can take a very general form, as illustrated in Figure 6 further below.
For the weight vector $w$ we choose the product of normal distributions as prior and independently the Gamma distribution for the standard deviation $\sigma$,
@@
This choice of prior imposes $\ell_{2}$-regularisation on $w$ (Ridge regression), cf. (4). We also tested a Laplacian prior (equivalent to $\ell_{1}$-regularisation) for $w$ but did not find any notable improvement. The conjugate prior for $w$ and $\sigma^{2}$ would be the (multivariate) normal-inverse-gamma distribution, but we prefer to set the priors for $w$ and $\sigma$ independently. This approach facilitates the empirical Bayes method used below to reduce the number of hyperparameters and has the additional benefit of not fixing the form of the posterior to the conjugate family, thus adding to the generality of the model.
The parameters $\mu_{w}$ and $\sigma_{w}$ of the prior for weights $w$ are determined in an empirical Bayes approach. For the mean we choose the ordinary least-square (OLS) estimate $\mu_{w}=C\,\bar{X}^{\mathrm{T}}\,\bar{y}$ with $C=(\bar{X}^{\mathrm{T}}\,\bar{X})^{-1}$. For the standard deviation we make use of the normality of OLS estimates and set $\sigma_{w}=\frac{1}{n}\,S\,\mathrm{diag}(C)$ with $S=(\bar{y}-\mu_{w}\,\bar{X}^{\mathrm{T}})\,\bar{y}$. Similarly, we set the mode $m_{\sigma}=\frac{\alpha-1}{\beta}$ of the Gamma distribution to $1$ to account for standardised data, i.e. $\alpha=m_{\sigma}\beta+1$. Finally, to still ensure weakly informative priors, we found that $\beta=\frac{1}{2}$ is an adequate choice, which corresponds to a variance of $\frac{\alpha}{\beta^{2}}=6$. Other moderate choices for the only remaining hyperparameter $\beta$ turned out to not have a significant influence on the results.
Figure 5: A schematic overview of the data model (top row) and the forecast model (bottom row). The data model takes the design matrix $X(d,h,\tau)$ and target values $y(d,h)$ as input, cf. (1) and (2). This input data is then organised according to forecast scenarios described in Section 5, where either the forecast creation time $\tau$ or the lag $h_{\mathrm{lag}}$ between forecast creation and delivery begin is fixed. Subsequently, the data is cleaned for missing values, standardised and undergoes feature selection (OMP for scenarios (a)-(e) and LASSO for (f)), as described in Section 3.4. The historical data $(\bar{X},\bar{y})$ enters the forecast model for learning, the new data point $(X^{\ast},y^{\ast})$ is used for forecast creation and evaluation. The hyperparameters $\mu_{w}$, $\sigma_{w}$, $\alpha$, $\beta$ of the prior $p_{\mathrm{prio}}$ are defined in an empirical Bayes approach using $(\bar{X},\bar{y})$, as explained in Section 4. The posterior $p_{\mathrm{post}}$ follows with Bayes formula from prior $p_{\mathrm{prio}}$ and likelihood $p_{\mathrm{li}}$, and the estimated posterior predictive distribution $\hat{p}_{\mathrm{ppd}}$ as the compound distribution of $p_{\mathrm{post}}$ and $p_{\mathrm{li}}$ at the new datapoint $X^{\ast}$, both explicated in Section 4. The procedure to extract point estimates $\hat{y}$ and prediction intervals $\hat{I}$ at credibility levels $\hat{\alpha}$ from $\hat{p}_{\mathrm{ppd}}$ is explained in Section 5.1, and their evaluation with respect to the true value $y^{\ast}$ and the live IDFull as benchmark is given in Section 5.2 in terms of mean absolute error (MAE), average coverage error (ACE), continuous ranked probability score (CRPS), and signs of spread and rest values.
@@
Bayesian point estimates for $w$ and $\sigma$ may now be obtained by maximizing the posterior $p_{\mathrm{post}}(w,\sigma\,|\,\bar{X},\bar{y})$ (MAP estimates), which plugged into the likelihood (5) and replacing $\bar{X}$ by $X^{\ast}$ yields a probabilistic forecast for the target $Y$. However, in doing so, we would discard valuable information about the uncertainty of $w$ and $\sigma$ captured by the posterior distribution. To fully incorporate this uncertainty, we instead determine the posterior predictive distribution (PPD),
@@
The data model and the forecast model are summarised in Figure 5.
@@ 5 Forecast study and discussion
Figure 6: A typical example of a forecast on the left (day $d=$ 24 August 2022, $h=0$) and a select example of a forecast on the right (day $d=$ 5 July 2022, hour $h=7$). The forecasts are represented by estimated posterior predictive distributions $\hat{p}_{\mathrm{ppd}}(y)$, from which credible intervals and point estimates are extracted. The example on the right, an extreme multimodal case showcasing the generality of the model, illustrates the procedure to pick a prediction interval and a point estimate using highest density intervals (HDIs). The intervals numbered 1, 2, 3, and 4 are formed by intersections with horizontal cuts, cf. (11). These intervals are special in that they would fuse into larger intervals for a slightly lower cut. From these fusing intervals, the interval with the largest credibility-to-width ratio is chosen, cf. (12), which in this example is interval 1. This interval serves as the PI $\hat{I}$, and the median within this interval as the point estimate $\hat{y}$, cf. (13).
@@
XScenario (a), 7 July 2022 XX Scenario (e), 5 October 2022
Figure 7: Exemplary full day forecasts for all delivery hours $h$ of day $d=\;$7 July 2022 on the left and all delivery hours $h$ of day $d=\;$5 October 2022 on the right. The forecast on the left is part of scenario (a), where the forecast creation time $\tau$ is fixed to 23:00 pm on day $d-1$, implying an increasing forecast lag of $h_{\mathrm{lag}}=h+1$ along the horizontal axis. The forecast on the right is part of scenario (e), where the lag between forecast and delivery begin is fixed to 1 hour, i.e. $h_{\mathrm{lag}}=1$, implying forecast creation times $\tau=h-1$. The vertical colour bars represent the topology of the respective predictive distributions $\hat{p}_{\mathrm{ppd}}(y)$ by means of $\alpha(p_{\mathrm{cut}})$ for 100 values of $p_{\mathrm{cut}}$, cf. (11).
@@ 5.1 Probabilistic forecasts
We illustrate our forecast procedure with a few examples. Two individual forecasts are shown in Figure 6. Since a typical reference price for the CID market is the DA price $P_{\mathrm{da}}$, we consider the spread between the IDFull and the DA price, $Y-P_{\mathrm{da}}$. Due to the probabilistic forecast, there are several options to extract a point estimate from the estimated predictive distribution $\hat{p}_{\mathrm{ppd}}(y)$, e.g. the mode, median or mean of $\hat{p}_{\mathrm{ppd}}(y)$. In terms of modes, an additional ambiguity arises when $\hat{p}_{\mathrm{ppd}}(y)$ is multimodal. Once a point estimate is extracted from $\hat{p}_{\mathrm{ppd}}(y)$, credible intervals of various credibility levels $\alpha$ can be extracted in various ways, which shall serve as prediction intervals (PIs) in this study. We found the following procedure to be best in terms of robustness and accuracy in our forecast study.
@@
For intermediate $p_{\mathrm{cut}}$, sub-intervals may develop which fuse again to larger credible intervals as $p_{\mathrm{cut}}$ is decreased. In order to choose a PI from all intervals $[l_{i}(p_{\mathrm{cut}}),u_{i}(p_{\mathrm{cut}})]$, we consider all fusing intervals $\{I_{\mathrm{fuse}}\}$, and of these pick $\hat{I}=[\hat{y}_{l},\hat{y}_{u}]$ that maximizes the credibility-to-width ratio,
@@
This procedure to determine PIs $\hat{I}$ and point estimates $\hat{y}$ in the general case of multimodal predictive distributions is also illustrated in Figure 6. In Figure 7, we show two examples in which the probabilistic forecasts for all hours of a day are depicted. These full day forecasts demonstrate the typical behaviour of our forecast model, namely identifying the live IDFull as the most important regressor, supporting the weak-form efficiency hypothesis. Nevertheless, the point estimates $\hat{y}$ often do correct the live IDFull values towards the ground truth $y^{\ast}$. We furthermore observe that forecasts with a larger lag between creation time $\tau$ and delivery begin $h$ are characterised by broader prediction intervals, indicating that our forecast model correctly represents forecast uncertainty.